# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

### v1

@@ Appendix B Performance Under Resource Constraints
Table 4: Performance of ZeroTuning Under Different Resource Constraints
@@
Appendix C Sensitivity to Different Context Lengths
@@
Table 5: Impact of Context Length on ZeroTuning Performance
@@
Appendix D Robustness Across Few-Shot Scenarios
Few-shot learning has become a widely adopted approach to improve the performance of LLMs by providing a small number of in-context examples, enabling adaptation to specific tasks with minimal data [52]. Building on previous zero-shot evaluations, we now evaluate the robustness of ZeroTuning in 1-shot and 2-shot scenarios across four datasets: SST-5, BoolQ, MMLU, and AQUA. To ensure consistency, we fix the randomly selected examples, maintain the selected head and scaling factor throughout the experiments.
@@ Appendix D Robustness Across Few-Shot Scenarios
Table 6: Comparison of Vanilla and ZeroTuning Performance Across Few-Shot Learning Scenarios
@@
Appendix E Impact of Decoding Strategies
@@
Appendix F The Effect of Different Numbers of Heads
@@
Appendix G Sensitivity to Prompt Variations
Prompts play a crucial role in guiding LLM behavior and typically consist of three components: Instruction1 (task guidance), Question (the actual query), and Instruction2 (output format specification). To evaluate the robustness of ZeroTuning under prompt perturbations, we perform experiments on the LLaMA-3.1-8B model using MMLU and SST-2 under three prompt formats: Full Prompt (Instruction1 + Question + Choices + Instruction2), Drop Instruction1, and Modify Instruction2. Detailed prompt examples are provided in Appendix J.
@@ Appendix G Sensitivity to Prompt Variations
Appendix H The Effect of Different Quantization Configurations
@@
Appendix I Limitations
Our current approach still relies on access to the ground-truth signals and accuracy to guide the tuning direction. As a result, it cannot yet operate in a fully unsupervised manner based solely on the properties of the input or the model’s attention distribution. Although we explored the effectiveness of the method under resource-constrained scenarios, such as using fixed scaling factors or skipping exhaustive traversal and classification of attention heads, these simplified variants consistently underperformed compared to the full method.
Additionally, our experiments follow standard evaluation protocols used in prior work [45, 46, 47, 12], and due to computational and inference resource constraints, we do not evaluate performance on longer-context generation tasks involving complex reasoning or multi-step thinking.
Future work could investigate the differences in attention distributions across layers and heads more deeply, which may lead to improved unsupervised tuning strategies. In addition, it may be fruitful to explore other tuning schemes, such as dynamically adjusting attention at critical positions in the long generated sequence.
Appendix J Prompts Used for Each Dataset
@@ Appendix I Limitations
For multiple choice task, we use the following prompt:
Prompt for Multiple-Choice Tasks Generate the correct answer to the following question. $<$Question$>$ $<$choice 1$>$ $<$choice 2$>$ $<$choice 3$>$ Answer:"
@@
Prompt for SST-2 "Classify the sentiment of the user’s message into one of the following categories: ’positive’ or ’negative’. Sentence: $<$sentence$>$ Sentiment: "
Prompt for SST-5 "Classify the sentiment of the user’s message into one of the following categories: ’terrible’, ’negative’, ’neutral’, ’positive’, or ’great’. Sentence: $<$sentence$>$ Sentiment: "
Prompt for MR "Classify the sentiment of the movie’s review into one of the following categories: ’positive’ or ’negative’. Review: $<$sentence$>$ Sentiment: "
Prompt for TREC "Classify the given questions into the following categories: ’Description’, ’Entity’, ’Expression’, ’Person’, ’Number’, or ’Location’. Question: $<$sentence$>$ Type: "
Prompt for CB "Read the following paragraph and determine if the hypothesis is true. Premise: $<$premise$>$ Hypothesis: $<$hypothesis$>$. Answer: "
Prompt for BoolQ "Read the text and answer the question by True or False. Text: $<$passage$>$ Question: $<$question$>$? Answer: "
Prompt for SUBJ "Classify the input into one of the following categories: subjective or objective. Input: $<$text$>$ Category: "
Appendix K Prompt for key tokens identification
@@
Appendix L LLM Output Example
L.1 Examples of SST-2 Dataset
Example 1:
Sentiment:
LLM Output: negative
Ground Truth: positive
Example 2:
Sentiment:
LLM Output: negative
Ground Truth: positive
Example 3:
Sentiment:
LLM Output: neutral
Ground Truth: positive
Example 4:
Answer:
LLM Output: The
Ground Truth: positive
L.2 Examples of BoolQ Dataset
Example 1:
Read the text and answer the question by True or False.
Text: Countdown (game show) – The contestant in control chooses six of 24 shuffled face-down number tiles, arranged into two groups: 20 “small numbers” (two each of 1 through 10), and four “large numbers” of 25, 50, 75 and 100. Some special episodes replace the large numbers with 12, 37, 62 and 87. The contestant decides how many large numbers are to be used, from none to all four, after which the six tiles are randomly drawn and placed on the board. A random three-digit target number is then generated by an electronic machine, affectionately known as “CECIL” (which stands for Countdown’s Electronic Calculator In Leeds). The contestants have 30 seconds to work out a sequence of calculations with the numbers whose final result is as close to the target number as possible. They may use only the four basic operations of addition, subtraction, multiplication and division, and do not have to use all six numbers. A number may not be used more times than it appears on the board. Fractions are not allowed, and only positive integers may be obtained as a result at any stage of the calculation. As in the letters rounds, any contestant who does not write down their calculations in time must go first, and both contestants must show their work to each other if their results and calculations are identical.
Question: do you have to use all the numbers on countdown?
Answer:
LLM Output: True
Ground Truth: False
Example 2:
Read the text and answer the question by True or False.
Question: will hawaii five o have a season 8?
Answer:
LLM Output: False
Ground Truth: True

### v2

@@ Appendix H Performance Under Resource Constraints
Table 5: Performance of ZeroTuning Under Different Resource Constraints.
@@
Appendix I Sensitivity to Different Context Lengths
@@
Table 6: Impact of Context Length on ZeroTuning Performance.
@@
Appendix J Robustness Across Few-Shot Scenarios
Few-shot learning has become a widely adopted approach to improve the performance of LLMs by providing a small number of in-context examples, enabling adaptation to specific tasks with minimal data (Brown et al., 2020). Building on previous zero-shot evaluations, we now evaluate the robustness of ZeroTuning in 1-shot and 2-shot scenarios across four datasets: SST-5, BoolQ, MMLU, and AQUA. To ensure consistency, we fix the randomly selected examples, maintain the selected head and scaling factor throughout the experiments.
@@
Table 7: Comparison of Vanilla and ZeroTuning Performance Across Few-Shot Learning Scenarios.
@@
Appendix K Impact of Decoding Strategies
@@
Appendix L The Effect of Different Numbers of Heads
@@
Appendix M Sensitivity to Prompt Variations
Prompts play a crucial role in guiding LLM behavior and typically consist of three components: Instruction1 (task guidance), Question (the actual query), and Instruction2 (output format specification). To evaluate the robustness of ZeroTuning under prompt perturbations, we perform experiments on the LLaMA-3.1-8B model using MMLU and SST-2 under three prompt formats: Full Prompt (Instruction1 + Question + Choices + Instruction2), Drop Instruction1, and Modify Instruction2. Detailed prompt examples are provided in Appendix O.
@@
Appendix N The Effect of Different Quantization Configurations
@@
Appendix O Prompts Used for Each Dataset
@@
For multiple-choice task, we use the following prompt:
Prompt for Multiple-Choice Tasks Generate the correct answer to the following question. <<Question>> <<choice 1>> <<choice 2>> <<choice 3>> Answer:”
@@
Prompt for SST-2 ”Classify the sentiment of the user’s message into one of the following categories: ’positive’ or ’negative’. Sentence: <<sentence>> Sentiment: ”
Prompt for SST-5 ”Classify the sentiment of the user’s message into one of the following categories: ’terrible’, ’negative’, ’neutral’, ’positive’, or ’great’. Sentence: <<sentence>> Sentiment: ”
Prompt for MR ”Classify the sentiment of the movie’s review into one of the following categories: ’positive’ or ’negative’. Review: <<sentence>> Sentiment: ”
Prompt for TREC ”Classify the given questions into the following categories: ’Description’, ’Entity’, ’Expression’, ’Person’, ’Number’, or ’Location’. Question: <<sentence>> Type: ”
Prompt for CB ”Read the following paragraph and determine if the hypothesis is true. Premise: <<premise>> Hypothesis: <<hypothesis>>. Answer: ”
Prompt for BoolQ ”Read the text and answer the question by True or False. Text: <<passage>> Question: <<question>>? Answer: ”
Prompt for SUBJ ”Classify the input into one of the following categories: subjective or objective. Input: <<text>> Category: ”
Appendix P Prompt for key tokens identification